Two particles having equal mass and carrying charg
Answer #1

A) potential by charge Q is V =kQ/(r)

Where k=coulomb constant

B) work needed w = q*V =k*Q*2Q/r

C) from conservation of momentum

Intial total momentum =final total momentum

=> 0= m*v+m*V

=> v & V are equal in magnitude and opposite in direction

From conservation of energy

Electric P.E = K.E1+K.E2

=> k*Q*2Q/r = 2*1/2*m*v^2

=>v^2=2*k*Q^2/m =>v=sqrt (2*k*Q^2/m)